the organizers of a charity concert sold tickets at two different prices. if they had sold 112 of the dearer tickets and 60 of the cheaper ones they would have received $45.60, but if they had sold 96 of the dearer tickets and 120 of the cheaper ones they would have received $52.80. find the price of the dearer tickets.
Answer by mananth(16946)

x= cost of dearer ticket
y= cost of cheaper ticket

112.00 x + 60.00 y = 45.60 .............1
Total value
96.00 x + 120.00 y = 52.80 .............2
Eliminate y
multiply (1)by -2.00
Multiply (2) by 1.00
-224.00 x -120.00 y = -91.20
96.00 x + 120.00 y = 52.80
Add the two equations
-128.00 x = -38.40
/ -128.00
x = 0.30
plug value of x in (1)
112.00 x + 60.00 y = 45.60
33.60 + 60.00 y = 45.60
60.00 y = 45.60 -33.60
60.00 y = 12.00
y = 0.20
x= 0.30 cost of dearer ticket
y= 0.20 cost of cheaper ticket
m.ananth@hotmail.ca